Find the laplacian of the function (phi(x,y,z)=2x^2y-xzn^3
1
Expert's answer
2019-10-16T13:21:57-0400

We are given "\phi(x,y,z,) = 2 x^2 y - x z n^3".

The Laplacian of "\phi(x,y,z)" in Euclidean coordinates is "\triangle \phi(x,y,z) = \frac{\partial^2 \phi}{\partial x^2} + \frac{\partial^2 \phi}{\partial y^2} + \frac{\partial^2 \phi}{\partial z^2}". Let us calculate the derivatives:

"\frac{\partial \phi}{\partial x} = 4 x y - z n^3", "\frac{\partial^2 \phi}{\partial x^2} = 4 y",

"\frac{\partial \phi}{\partial y} = 2 x^2", "\frac{\partial^2 \phi}{\partial y^2} = 0",

"\frac{\partial \phi}{\partial z} = - x n^3", "\frac{\partial^2 \phi}{\partial z^2} = 0".

Hence, "\triangle \phi(x,y,z) = 4 y".